OpenStudy (anonymous):

How to Write: Log2 (Nt-N0 ) As a logarithm ln:

OpenStudy (anonymous):

Or Log2 (A-B) as log Ln

OpenStudy (anonymous):

\[\frac{ Log _{e}(Nt-N0) }{ Log _{e}2 }\] Something like this e = Neper number

OpenStudy (anonymous):

\[\ln (Nt-No) / \ln 2\]

OpenStudy (anonymous):

Good @duck123